# thanks: deceasedthanks: deceased

### v1

@@ III.2 Selection of final state
For all the selected events, $[pp\pi^{-}\pi^{-}]+p_{\rm miss}$ and $[pn\pi^{+}\pi^{-}]+n_{\rm miss}$, we applied a common procedure to identify the final states: $[\pi^{-}\Lambda p]+p_{\rm miss}$ (Eq.(3)), $[\pi^{+}\Lambda n]+n_{\rm miss}$ (Eq.(4)), and $[\pi^{\mp}\Sigma^{\pm}p]+n_{\rm miss}$ (Eq.(5)). We used a log-likelihood method, which was introduced in Ref. [2], to determine whether or not the event kinematics and topology were consistent with the reactions described in Eqs.(3), (4), and (5). To assess the kinematic information, we performed a kinematic fit to evaluate how well the event kinematics matched the final state of interest within the detector’s resolution. For geometrical information, we utilized the distances of the closest approach (DCA) between two charged particles. These distances should be zero within the resolution of the detector.
@@
where $N_{\text{DCA}}$ is the number of DCAs considered. $p(\chi^{2}_{\text{kfit}})$ represents the probability function of the $\chi^{2}$ with ${\text{NDF}}_{\text{kfit}}=2$. $p(\text{DCA})$ is a Gaussian function with mean $\mu=0$ and standard deviation $\sigma$, evaluated through Monte Carlo simulation with realistic detector resolutions. All $p(x)$ functions are normalized such that their maximum values are equal to one, making $L=0$ at the most probable point.
Multiple-particle combinations could be generated by hyperon decays or $\bar{K}^{0}\to\pi^{+}\pi^{-}$ decay. We examined all $L$ with possible particle combinations for each event to isolate which particle pairs are the decay products. Then, we determined that the reaction process of the events resulted in the minimum $L$. Figure 2 illustrates the distributions of the $L/{\rm NDF}$ for each identified process. As shown in the figure, the $L/{\rm NDF}$ distributions are quite similar, with the most frequent $L/{\rm NDF}$ values clustering around $L/{\rm NDF}\sim 0.5$. The yields decrease as $L/{\rm NDF}$ becomes larger. To select the final event sets, we accepted events with $L/{\rm NDF}<2$, and the threshold is indicated by the vertical red line in the figure.
@@
Due to the finite resolution, the correct particle combination for the reaction process is not always associated with the minimum $L$. There are three significant sources of contamination: I) combinatorial effects (improper particle pairing in hyperon decay, including $K^{0}$ contamination), II) $\Sigma^{0}$ contamination (resulting from a missing $\gamma$ in the $\Sigma^{0}$ decay), and III) contamination arising from the $\pi\pi Y\!N\!N$ final states. We assessed the contamination yields in each channel through simulation, as shown in the hatched region in the figure. The relative background yields to the total events in the $[\pi^{-}\Lambda p]+p_{\rm miss}$, $[\pi^{+}\Lambda n]+n_{\rm miss}$, and $[\pi^{\mp}\Sigma^{\pm}p]+n_{\rm miss}$ event sets were found to be $({\rm I}+{\rm II}+{\rm III})\sim(0+15+0)\%$, $(4+12+6)\%$, and $(5+0+6)\%$, respectively.
To demonstrate the validity of the event selection, we present the event distributions of the invariant mass of the selected $\pi N$ pairs from the hyperon decay ($m_{\pi N}$) and the missing mass of $\pi Y\!N$ ($m_{X}$) in Fig. 3. The top panel of the figure pair displays events with $L<100$, while the bottom panel represents events after the selection. The signals corresponding to each $[\pi Y\!N]+N_{\rm miss}$ final state are clearly observed as concentrated events around $(m_{X},~{}m_{\pi N})\sim(m_{N},~{}m_{Y})$, where $m_{N}$ and $m_{Y}$ represent the intrinsic masses of the nucleon and $\Lambda$ or $\Sigma$, respectively.
Figure 3: Two-dimensional distributions of the masses of the hyperon candidate ($m_{N\pi}$) and the missing particle ($m_{X}$). Panels (a), (b), (c), and (d) correspond to the event selections of $\pi^{-}\Lambda pp$, $\pi^{+}\Lambda nn$, $\pi^{-}\Sigma^{+}pn$, and $\pi^{+}\Sigma^{-}pn$, respectively. The top figures display all events with $L<100$, while the bottom figures show the selected events with $L<2~{}{\rm NDF}$.
@@ III.3 Acceptance correction
In the measurement, the acceptance of an event is determined by the four particles to be measured, $[\pi Y\!N]\to[\pi(\pi_{Y}N_{Y})N]$. We consider the event kinematics as a cascading two-body decay scheme illustrated in Fig. 4.
Figure 4: Considered kinematics for the $K^{-}+^{3}{\rm He}\to[\pi Y\!N]+N_{\rm miss}$ reaction. The figure illustrates the six kinematical parameters (indicated by red) that we considered for the acceptance correction.
@@
as shown in the figure. The kinematics of each vertex point are defined by three parameters: the polar angle ($\theta$) and azimuthal angle ($\phi$) of the daughters in the rest frame of the parent system and the parent mass. The masses of the initial $K^{-}+~{}^{3}{\rm He}$ system and the final $\pi_{Y}+N_{Y}$ system are constrained to be $\sqrt{s}$ and $m_{Y}$, respectively. Therefore, there are ten independent kinematical parameters: the masses of $[\pi Y\!N]$ and $[\pi Y]$, $\theta$ and $\phi$ of $N_{\rm miss}$, $N$, $\pi$, and $\pi_{Y}$.
Since the azimuthal angles ($\phi$) do not affect the kinematics of the successive process, we can accurately evaluate the acceptance by integrating out $\phi$, assuming a uniform distribution. However, the polar angles ($\theta$) cannot be integrated. Consequently, we must consider the six kinematic parameters in the figure for the acceptance correction.
@@
Figure 5: Distributions of the five kinematical parameters considered in the acceptance correction for the $[\pi^{-}\Lambda p]+p_{\rm miss}$ channel. The black and blue distributions represent the data, and the four-body phase space distribution, including the acceptance. The simulated distributions were scaled to have the same integral as the data.
If the reaction is governed by the phase space only without any additional physics effects beyond the density of states of the final state, we would expect the data (black) and the simulation (blue) to coincide.
As shown in Fig. 5 (a) to (c), the data deviates from the simulation, but there is a good agreement in (d) to (f). Hence, we applied the acceptance correction event-by-event basis for these three parameters while integrating and correcting the remaining two using the average acceptance. Furthermore, we used $q_{[\pi Y\!N]}$ instead of $\theta_{N_{\rm miss}}$ since our interest lies in the momentum transfer to the $[\pi Y\!N]$ system rather than the specific formation angle $\theta_{N_{\rm miss}}$. Consequently, we constructed a three-dimensional acceptance map, $\varepsilon(m_{[\pi Y\!N]},~{}q_{[\pi Y\!N]},~{}m_{[\pi Y]})$, which encompasses the geometric acceptance of the CDS, detection efficiencies, and analysis efficiencies.
We are specifically interested in three distributions, namely $m_{[\pi Y\!N]}$, $q_{[\pi Y\!N]}$, and $m_{[\pi Y]}$, for each $[\pi Y\!N]+N_{\rm miss}$ channel to investigate the production of $\bar{K}\!N\!N$ or $Y^{*}$ resonances. To obtain the differential cross sections of these three parameters, we initially calculated the triple differential cross section $d^{3}\sigma/dm\,dq\,dm_{[\pi Y]}$, where $m$ represents $m_{[\pi Y\!N]}$, $q$ represents $q_{[\pi Y\!N]}$. The expression is given by:
@@
where $\mathcal{L}=2.89\pm 0.01~{}{\rm nb}^{-1}$ represents the integrated luminosity, $N(m,\,q,\,m_{[\pi Y]})$ is the number of events in the ($m,\,q,\,m_{[\pi Y]}$) bin with bin widths $\Delta m$, $\Delta q$, and $\Delta m_{[\pi Y]}$. $N_{\rm fake}(m,\,q,\,m_{[\pi Y]})$ denotes the expected fake-neutron background yield in the real data, as described in Sec. III.2 (note that $N_{\rm fake}(m,\,q,\,m_{[\pi Y]})=0$ for the $[\pi^{-}\Lambda p]+p_{\rm miss}$ channel, since there are no neutrons in this channel).
We omitted low-efficiency region $\varepsilon(m,q,m_{[\pi Y]})<0.1\%$ for the $[\pi^{-}\Lambda p]+p_{\rm miss}$, and $\varepsilon(m,q,m_{[\pi Y]})<0.015\%$ for other channels. Finally, we obtained differential cross-sections of interest by integrating the triple-differential cross-sections.
@@ IV Results
Figure 6: Acceptance-corrected $(m_{[\pi Y\!N]},q_{[\pi Y\!N]})$ distributions and their projections on the $m_{[\pi Y\!N]}$ axes are shown in the figure: (a) corresponds to $[\pi^{-}\Lambda p]+p_{\rm miss}$, (b) corresponds to $[\pi^{+}\Lambda n]+n_{\rm miss}$, (c) corresponds to $[\pi^{-}\Sigma^{+}p]+n_{\rm miss}$, and (d) corresponds to $[\pi^{+}\Sigma^{-}p]+n_{\rm miss}$. The events-concentrations along a kinematical line (Eq.(10)) are clearly shown, especially in (a). For the mass region $m_{\pi^{-}\Lambda p}>2.6~{}{\rm GeV}/c^{2}$, the cross-section in (a) is scaled by a factor of $1/5$ to show events-concentration corresponding to Eq.(11). The black dotted lines represent the kinematical limit of the reaction. The blue vertical dotted lines indicate the mass threshold of $m_{\bar{K}}+2m_{N}$. The red dotted curves represent the quasi-free kinematical line described in Eq.(10).
@@
An events-concentration at $(m_{[\pi Y\!N]},~{}q_{[\pi Y\!N]})\sim(2.4~{}{\rm GeV}/c^{2},~{}0.3~{}{\rm GeV}/c)$ was observed consistently in all $\pi Y\!N\!N$ channels. The origin of these events-concentrations can be understood as follows. The concentrations are along the kinematical line of the quasi-free $\bar{K}$ absorption by residual two nucleons given by
@@
where $m_{N}$ and $m_{\bar{K}}$ are the intrinsic masses of the $N$ and $\bar{K}$, respectively. Equation (10) describes the invariant mass of $\bar{K}$ with momentum $q$ interacting with two residual nucleons, thereby representing the kinematics of a quasi-free absorption process where two residual nucleons absorb an on-shell $\bar{K}$ at rest. Given that the events are concentrated along this kinematical line, we can interpret this concentration as a manifestation of the quasi-free $\bar{K}$ absorption (${\rm QF}_{\bar{K}-{\rm abs}}$) process, similar to the non-mesonic $[\Lambda p]+n_{\rm miss}$ channel [2].
If the $\bar{K}\!N\!N$ quasi-bound state observed in the $[\Lambda p]+n_{\rm miss}$ channel [2] decay into the mesonic channels, one can expect to observe spectral peaks in the invariant-mass spectra of each $[\pi Y\!N]$ channel below the kaon binding threshold, $m_{\bar{K}\!N\!N}=m_{\bar{K}}+2m_{N}$ (indicated by the blue dotted line in the figure). However, clear peak structures were not found in these spectra. The absence of clear peak structures can be understood by considering the density of states in the final state, which is governed by the four-body phase space of $[\pi Y\!N]+N_{\rm miss}$. The phase space volume below the $m_{\bar{K}\!N\!N}$ threshold becomes very small since the mass region is close to the $\pi Y\!N$ threshold, and the four-body phase space opens from the $\pi Y\!N$ threshold only in the second order. As a result, the invariant-mass spectra of $[\pi Y\!N]$ in this region are significantly modified due to the limited phase space near the threshold. This spectral modification leads to the disappearance of the prominent peak structure. The density of states in the mesonic final states and the resulting spectral modification will be described in more detail in Sec. V.
@@
In this process, the $\bar{K}$ reacts only with a deuteron-like momentum correlated $pn$-cluster in ${{}^{3}\rm He}$. The other proton in ${}^{3}{\rm He}$ acts as a spectator with a Fermi momentum, serving as the missing particle ($p_{\rm miss}$). We plotted the momentum distribution of $p_{\rm miss}$ in Fig. 7 to investigate this interpretation. As shown in the figure, the momentum distribution of $p_{\rm miss}$ is consistent with our spectator picture, exhibiting a low momentum component in agreement with Fermi motion in ${{}^{3}\rm He}$ [15].
Figure 7: Momentum distribution of the missing proton ($p_{\rm miss}$) in the $(\pi^{-}\Lambda p+p_{\rm miss})$ channel. The gray dotted and solid lines represent the Fermi momentum distribution in ${{}^{3}\rm He}$ [15].
In contrast, the same events-concentration was not observed in the $[\pi^{+}\Lambda n]+n_{\rm miss}$ channels, as shown in Fig. 6-(b). This result suggests that the $K^{-}$ beam is less likely to be absorbed by a ($pp$) pair compared to a ($pn$) pair, whose ratio ($pp$/$pn$) is less than $1/10$. This result is consistent with the absorption of stopped-$\pi^{-}$ on ${{}^{3}\rm He}~{}$[16].

### v2

@@ III.2 Selection of final state
For all the selected events, $[pp\pi^{-}\pi^{-}]+p_{\rm miss}$ and $[pn\pi^{+}\pi^{-}]+n_{\rm miss}$, we applied a common procedure to identify the final states: $[\pi^{-}\Lambda p]+p_{\rm miss}$ (Eq. (3)), $[\pi^{+}\Lambda n]+n_{\rm miss}$ (Eq. (4)), and $[\pi^{\mp}\Sigma^{\pm}p]+n_{\rm miss}$ (Eq. (5)). We used a log-likelihood method, which was introduced in Ref. [2], to determine whether the event kinematics and topology are consistent with the reactions described in Eqs. (3), (4), and (5). To assess the kinematic information, we conducted a kinematic fit to evaluate how well the event kinematics matched the final state of interest within the detector’s resolution. For geometrical information, we used the distances of the closest approach (DCA) between two charged particles. These distances should be zero within the resolution of the detector.
@@
where $N_{\text{DCA}}$ is the number of DCAs considered and $p(\chi^{2}_{\text{kfit}})$ represents the probability function for $\chi^{2}$ with ${\text{NDF}}_{\text{kfit}}=2$. Function $p(\text{DCA})$ is a Gaussian function with mean $\mu=0$ and standard deviation $\sigma$, evaluated through Monte Carlo simulation with realistic detector resolutions. All of the $p(x)$ functions are normalized such that their maximum values are equal to one, making $L=0$ at the most probable point.
Multiple-particle combinations could be generated by hyperon decays or $\bar{K}^{0}\to\pi^{+}\pi^{-}$ decay. We examined all $L$ with possible particle combinations for each event to isolate which particle pairs are the decay products. We then determined that the reaction process of the events resulted in the minimum $L$. Figure 2 illustrates the distributions of the $L/{\rm NDF}$ for each identified process. As shown in the figure, the $L/{\rm NDF}$ distributions are quite similar, with the most frequent $L/{\rm NDF}$ values clustering around $L/{\rm NDF}\sim 0.5$. The yields decrease as $L/{\rm NDF}$ becomes larger. To select the final event sets, we accepted events with $L/{\rm NDF}<2$; the threshold is indicated by the vertical red line in the figure.
@@
Because of the finite resolution, the correct particle combination for the reaction process is not always associated with the minimum $L$. There are three significant sources of contamination: I) combinatorial effects (improper particle pairing in hyperon decay, including $K^{0}$ contamination), II) $\Sigma^{0}$ contamination (resulting from a missing $\gamma$ in the $\Sigma^{0}$ decay), and III) contamination arising from the $\pi\pi Y\!N\!N$ final states. We assessed the contamination yields in each channel through simulation, as shown in the hatched region in the figure. The background yields relative to the total events in the $[\pi^{-}\Lambda p]+p_{\rm miss}$, $[\pi^{+}\Lambda n]+n_{\rm miss}$, and $[\pi^{\mp}\Sigma^{\pm}p]+n_{\rm miss}$ event sets were found to be $({\rm I}+{\rm II}+{\rm III})\sim(0+15+0)\%$, $(4+12+6)\%$, and $(5+0+6)\%$, respectively.
To demonstrate the validity of the event selection, we present the event distributions of the invariant mass of the selected $\pi N$ pairs from the hyperon decay ($m_{\pi N}$) and the missing mass of $\pi Y\!N$ ($m_{X}$) in Fig. 3. The top panel of the figure pair displays events with $L<100$, whereas the bottom panel represents events after the selection. The signals corresponding to each $[\pi Y\!N]+N_{\rm miss}$ final state are clearly observed as concentrated events around $(m_{X},~{}m_{\pi N})\sim(m_{N},~{}m_{Y})$, where $m_{N}$ and $m_{Y}$ represent the intrinsic masses of the nucleon and $\Lambda$ or $\Sigma$, respectively.
Figure 3: Two-dimensional distributions of masses of hyperon candidate ($m_{N\pi}$) and missing particle ($m_{X}$). Panels (a), (b), (c), and (d) correspond to the event selections of $\pi^{-}\Lambda pp$, $\pi^{+}\Lambda nn$, $\pi^{-}\Sigma^{+}pn$, and $\pi^{+}\Sigma^{-}pn$, respectively. The top figures display all events with $L<100$, whereas the bottom figures show the selected events with $L<2~{}{\rm NDF}$.
@@ III.3 Acceptance correction
In the measurement, the acceptance of an event is determined by the four particles to be measured, $[\pi Y\!N]\to[\pi(\pi_{Y}N_{Y})N]$. We consider the event kinematics as a cascading two-body decay scheme as illustrated in Fig. 4.
Figure 4: Kinematics considered for $K^{-}+{{}^{3}{\rm He}}\to[\pi Y\!N]+N_{\rm miss}$ reaction. The figure illustrates the six kinematical parameters (indicated by red) considered for acceptance correction.
@@
as shown in the figure. The kinematics of each vertex point are defined by three parameters: the polar angle ($\theta$) and azimuthal angle ($\phi$) of the daughters in the rest frame of the parent system, and the parent mass. The masses of the initial $K^{-}+~{}^{3}{\rm He}$ system and the final $\pi_{Y}+N_{Y}$ system are constrained to be $\sqrt{s}$ and $m_{Y}$, respectively. Therefore, there are ten independent kinematical parameters: the masses of $[\pi Y\!N]$ and $[\pi Y]$; $\theta$ and $\phi$ of $N_{\rm miss}$; $N$; $\pi$; and $\pi_{Y}$.
Because the azimuthal angles ($\phi$) do not affect the kinematics of the successive process, we can accurately evaluate the acceptance by integrating $\phi$ under the assumption of a uniform distribution. However, the polar angles ($\theta$) cannot be integrated. Consequently, we must consider the six kinematic parameters in the figure for the acceptance correction.
@@
Figure 5: Distributions of the five kinematical parameters considered in the acceptance correction for the $[\pi^{-}\Lambda p]+p_{\rm miss}$ channel. The black and blue distributions represent the data and the four-body phase space distribution, including the acceptance. The simulated distributions were scaled to have the same integral as the data.
If the reaction is governed by the phase space only, without any additional physics effects beyond the density of states of the final state, we would expect the data (black) and the simulation (blue) to coincide.
As shown in Fig. 5(a)–(c), the data deviate from the simulation; however, good agreement is observed in Fig. 5(d)–(f). Hence, we applied the acceptance correction on an event-by-event basis for these three parameters while integrating and correcting the remaining three using the average acceptance. In addition, we used $q_{[\pi Y\!N]}$ instead of $\theta_{N_{\rm miss}}$ because our interest lies in the momentum transfer to the $[\pi Y\!N]$ system rather than in the specific formation angle $\theta_{N_{\rm miss}}$. Consequently, we constructed a three-dimensional acceptance map, $\varepsilon(m_{[\pi Y\!N]},~{}q_{[\pi Y\!N]},~{}m_{[\pi Y]})$, which encompasses the geometric acceptance of the CDS, detection efficiencies, and analysis efficiencies.
To investigate the production of $\bar{K}\!N\!N$ or $Y^{*}$ resonances, we are specifically interested in three distributions—specifically, $m_{[\pi Y\!N]}$, $q_{[\pi Y\!N]}$, and $m_{[\pi Y]}$—for each $[\pi Y\!N]+N_{\rm miss}$ channel. To obtain the differential cross-sections of these three parameters, we initially calculated the triple differential cross-section $d^{3}\sigma/dm\,dq\,dm_{[\pi Y]}$, where $m$ represents $m_{[\pi Y\!N]}$ and $q$ represents $q_{[\pi Y\!N]}$. The expression is given by
@@
where $\mathcal{L}=2.89\pm 0.01~{}{\rm nb}^{-1}$ represents the integrated luminosity and $N(m,\,q,\,m_{[\pi Y]})$ is the number of events in the ($m,\,q,\,m_{[\pi Y]}$) bin with bin widths $\Delta m$, $\Delta q$, and $\Delta m_{[\pi Y]}$, respectively. Parameter $N_{\rm fake}(m,\,q,\,m_{[\pi Y]})$ denotes the expected fake-neutron background yield in the real data, as described in Sec. III.2 (note that $N_{\rm fake}(m,\,q,\,m_{[\pi Y]})=0$ for the $[\pi^{-}\Lambda p]+p_{\rm miss}$ channel because this channel contains no neutrons).
We omitted the low-efficiency region $\varepsilon(m,q,m_{[\pi Y]})<0.1\%$ for the $[\pi^{-}\Lambda p]+p_{\rm miss}$ and the region $\varepsilon(m,q,m_{[\pi Y]})<0.015\%$ for the other channels. Finally, we obtained differential cross-sections of interest by integrating the triple-differential cross-sections.
@@ IV Results
Figure 6: Acceptance-corrected $(m_{[\pi Y\!N]},q_{[\pi Y\!N]})$ distributions and their projections on the $m_{[\pi Y\!N]}$ axes are shown in the figure: (a) corresponds to $[\pi^{-}\Lambda p]+p_{\rm miss}$, (b) corresponds to $[\pi^{+}\Lambda n]+n_{\rm miss}$, (c) corresponds to $[\pi^{-}\Sigma^{+}p]+n_{\rm miss}$, and (d) corresponds to $[\pi^{+}\Sigma^{-}p]+n_{\rm miss}$. The event concentrations along a kinematical line (Eq. (10)) are clearly shown, especially in (a). For the mass region $m_{\pi^{-}\Lambda p}>2.6~{}{\rm GeV}/c^{2}$, the cross-section in (a) is scaled by a factor of $1/5$ to show the events-concentration corresponding to Eq. (11). The black dotted lines represent the kinematical limit of the reaction. The blue vertical dotted lines indicate the mass threshold of $m_{\bar{K}}+2m_{N}$. The red dotted curves represent the quasi-free kinematical line described in Eq. (10).
@@
An event concentration at $(m_{[\pi Y\!N]},~{}q_{[\pi Y\!N]})\sim(2.4~{}{\rm GeV}/c^{2},~{}0.3~{}{\rm GeV}/c)$ was observed consistently in all of the $\pi Y\!N\!N$ channels. The origin of these event concentrations can be understood as follows. The concentrations are along the kinematical line of the quasi-free $\bar{K}$ absorption by two residual nucleons, given by
@@
where $m_{N}$ and $m_{\bar{K}}$ are the intrinsic masses of the $N$ and $\bar{K}$, respectively. Equation (10) describes the invariant mass of $\bar{K}$ with momentum $q$ interacting with two residual nucleons, thereby representing the kinematics of a quasi-free absorption process in which two residual nucleons absorb an on-shell $\bar{K}$ at rest. Given that the events are concentrated along this kinematical line, we can interpret this concentration as a manifestation of the quasi-free $\bar{K}$ absorption (${\rm QF}_{\bar{K}-{\rm abs}}$) process, similar to the non-mesonic $[\Lambda p]+n_{\rm miss}$ channel [2].
If the $\bar{K}\!N\!N$ quasi-bound state observed in the $[\Lambda p]+n_{\rm miss}$ channel [2] decays into the mesonic channels, spectral peaks can be observed in the invariant-mass spectra of each $[\pi Y\!N]$ channel below the kaon binding threshold, $m_{\bar{K}\!N\!N}=m_{\bar{K}}+2m_{N}$ (indicated by the blue dotted line in the figure). However, clear peak structures were not observed in these spectra. The absence of clear peak structures can be understood by considering the density of states in the final state, which is governed by the four-body phase space of $[\pi Y\!N]+N_{\rm miss}$. The phase space volume below the $m_{\bar{K}\!N\!N}$ threshold becomes small because the mass region is close to the $\pi Y\!N$ threshold, and the four-body phase space opens from the $\pi Y\!N$ threshold only in the second order. As a result, the invariant-mass spectra of $[\pi Y\!N]$ in this region are substantially modified because of the limited phase space near the threshold. This spectral modification leads to the disappearance of the prominent peak structure. The density of states in the mesonic final states and the resultant spectral modification are described in greater detail in Sec. V.
@@
In this process, the $\bar{K}$ reacts only with a deuteron-like momentum-correlated $pn$ cluster in ${{}^{3}\rm He}$. The other proton in ${}^{3}{\rm He}$ acts as a spectator with a Fermi momentum, serving as the missing particle ($p_{\rm miss}$). We plotted the momentum distribution of $p_{\rm miss}$ in Fig. 7 to investigate this interpretation. As shown in the figure, the momentum distribution of $p_{\rm miss}$ is consistent with our spectator picture, showing a low momentum component, in agreement with the Fermi motion in ${{}^{3}\rm He}$ [15].
Figure 7: Momentum distribution of missing proton ($p_{\rm miss}$) in $(\pi^{-}\Lambda p+p_{\rm miss})$ channel. The gray dotted and solid lines represent the Fermi momentum distribution in ${{}^{3}\rm He}$ [15].
By contrast, the same event concentration was not observed in the $[\pi^{+}\Lambda n]+n_{\rm miss}$ channels (Fig. 6(b)). This result suggests that the $K^{-}$ beam is less likely to be absorbed by a $pp$ pair than by a $pn$ pair, whose ratio ($pp$/$pn$) is less than $1/10$. This result is consistent with the absorption of stopped-$\pi^{-}$ on ${{}^{3}\rm He}~{}$[16].